For the equation (a-b)^2 + a^2 = 25, a and b are integers and a is greater than or equal to 0. List the ordered pairs that occur.
Answer by ikleyn(52908)

.

Recall that  3^2 + 4^2 = 25  is the only Pythagorean triple with  c^2 = 25.


It gives the list of pairs (a,b) as follows


    (a,b) = (3,-1), (4,1), (3,7), (4,7).

Answered and solved.